### v1

# PROTOCOL FOR AN OBSERVATIONAL STUDY ON THE EFFECTS OF COMBINATIONS OF ADVERSE CHILDHOOD EXPERIENCES ON ADULT DEPRESSION

Abstract.
Adverse childhood experiences (ACEs) have been linked to a wide range of negative health outcomes in adulthood. However, few studies have investigated what specific combinations of ACEs most substantially impact mental health. In this article, we provide the protocol for our observational study of the effects of combinations of ACEs on adult depression. We use data from the 2023 Behavioral Risk Factor Surveillance System (BRFSS) to assess these effects. We will evaluate the replicability of our findings by splitting the sample into two discrete subpopulations of individuals. We employ data turnover for this analysis, enabling a single team of statisticians and domain experts to collaboratively evaluate the strength of evidence, and also integrating both qualitative and quantitative insights from exploratory data analysis. We outline our analysis plan using this method and conclude with a brief discussion of several specifics for our study.
Corresponding authors:
William Bekerman, bekerman@wharton.upenn.edu; Ruizhe Zhang, zhangrz22@m.fudan.edu.cn
1. Introduction
Adverse childhood experiences (ACEs) describe specific categories of early life hardship or harm that occur before an individual reaches adulthood. Although these can take various forms, they are typically categorized into ten types involving abuse (physical, emotional, or sexual) and neglect (physical or emotional), as well as household challenges including mental illness, substance abuse, intimate partner violence, parental absence due to divorce or separation, or incarceration. Initially investigated in a landmark study conducted by the Centers for Disease Control and Prevention (CDC) and Kaiser Permanente, ACEs are highly prevalent and exhibit a strong dose-response relationship with numerous negative health and social outcomes in adulthood (Felitti et al. [1998]; Dube et al. [2003]; Anda et al. [2006]). Exposure to individual ACEs has been repeatedly associated with a range of adverse physical and emotional health consequences later in life. One study indicated that childhood emotional abuse increases the risk of lifetime depressive disorders, with adjusted odds ratios (ORs) of 2.7 for women and 2.5 for men (Chapman et al. [2004]). Depression also appears to be closely related to having experienced parental alcohol abuse (Anda et al. [2002]). Moreover, physical and verbal abuse have been found to be strongly associated with body weight issues and obesity (Williamson et al. [2002]).
Unfortunately, when children are exposed to one type of adversity, there is a high likelihood of exposure to additional ACEs (Dong et al. [2004]; Felitti [2002]). Numerous large-scale investigations have confirmed that these detrimental experiences can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al. [1998]; Chapman et al. [2004]). Accordingly, previous literature has examined closely the number of categories of ACEs, often referred to as ACE score. For instance, one study using data from 14 states revealed substantial variation across states in the percentage change in health outcomes for those exposed to at least four ACEs compared to those with fewer than four ACEs (Waehrer et al. [2020]). Physically, individuals with four or more ACEs face elevated risks of heart disease, liver disease, kidney disease, asthma, Alzheimer’s disease, stroke, metabolic disorders, obesity, diabetes, and certain cancers (for Disease Control and Prevention [2017]; for Youth Wellness [2014]; Merrick [2019]; Hughes et al. [2017]; Waehrer et al. [2020]). On the emotional and psychological side, ACE score strongly correlates with mental health issues such as depression, anxiety, and suicidal ideation (Hillis et al. [2004]; Chapman et al. [2004]). Meta-analyses further reinforce that an accumulation of at least 4 ACEs correlates with higher risks of chronic disease, poor mental health, and negative social outcomes, demonstrating the need for strategies to prevent ACEs and lessen their long-term repercussions (Hughes et al. [2017]).
A recent study challenged the overall ACE score and advised to break down the measure back into specific items (Fitzgerald and Bishop [2024]). Gaps also remain in our understanding of how the accumulation of ACEs influences health. We are therefore interested in which categories of ACEs—and which specific combinations of ACEs (e.g., physical abuse + sexual abuse, or physical abuse + emotional neglect + substance abuse)—are most harmful and warrant primary care. We focus our analysis on depression, one of the most prevalent and burdensome mental health conditions globally. Consequently, our focus is to evaluate the effect of different combinations of specific ACE items on depression and identifying a higher-risk group of sets of ACEs.
We will examine the impacts of combinations of ACEs on adult depression using data from the Behavioral Risk Factor Surveillance System (BRFSS) 2023, an ongoing health-related telephone survey system administered by the CDC with a random sample of over 400,000 individuals across the United States in 2023. This data includes detailed information from ten states that administered the CDC’s Adverse Childhood Experiences module in 2023. We can obtain all ten ACE categories proposed in the CDC-Kaiser study (Felitti et al. [1998]), and there are over 50,000 individuals with complete records of all ACE items. Among them, around 17,000 reported zero ACEs while the rest were exposed to at least one ACE. This large sample size will help facilitate powerful data analysis and explore the influencing mechanism of different ACEs combinations. The BRFSS also collected measurements of variables that may confound the relationship between exposure to ACEs and depression, such as those related to personal demographics. We will utilize those variables to form separate strata for treated and control individuals, enabling us to compare the outcomes of similar individuals in hopes of ameliorating confounding biases.
We employ a novel statistical design called data turnover (Bekerman et al. in preparation, see Bekerman et al. [2024a] for a description) to conduct this analysis. By integrating both qualitative and quantitative insights from data exploration, data turnover enables a single team of statisticians and domain experts to evaluate evidence from multiple data splits. Exploratory data analysis (EDA) allows us to inspect the data before making potentially questionable assumptions, detect anomalies or outliers that do not represent natural variation in the data, and formulate new hypotheses. It also helps identify variables that may not measure what we initially intended, indicating the need for a better measure or refined, data-informed hypotheses (Bekerman et al. [2024b]). Beyond exploration, replication is critical in observational studies for reinforcing the credibility of findings and sustaining trust in scientific knowledge. Previous work has combined some—but not all—aspects of exploration and replication within the same observational study. For instance, some methods allow for exploration without replication (e.g., Cox [1975]; Heller et al. [2009]), whereas others enable replication without exploration (e.g., Zhao et al. [2018]; Karmakar et al. [2019]). There is also research that supports both exploration and replication but requires two independent investigative teams (Roy et al. [2022]). In contrast, data turnover makes it possible to conduct both exploration and replication using a single research group.
In observational studies like ours, where treatment assignment is not under our direct control, obtaining consistent findings across multiple groups that may have different assignment mechanisms reinforces the evidence that the treatment itself is truly responsible for the observed effects (Rosenbaum [2001]; Rosenbaum [2015]). By analyzing data from two distinct subpopulations of BRFSS 2023 participants—residing in states that voted for the Democratic candidate in the 2024 US presidential election (”blue states”) and those residing in states that voted for the Republican candidate (”red states”) —we can assess the replicability of our findings. We split our data into blue states and red states, recognizing that these political orientations often shape distinct social, economic, and health care policies. A study focusing on population in East and West Germany has shown that socio-political context is a determinant of childhood maltreatment (Ulke et al. [2021]). In our context, for example, blue states may allocate more resources toward public health programs (e.g., adopt medicaid expansion) or have broader social safety nets, whereas red states may emphasize different strategies for healthcare funding and child protective services. Such policy and cultural variations can influence both the prevalence of ACEs and the mechanisms by which individuals are exposed to them, potentially resulting in different patterns of ACEs across these two groups. Demonstrating consistent results in both blue and red states thus enhances confidence that our findings are robust across varied socio-political contexts. By employing data turnover, we can predetermine how one subgroup will guide analysis of the other, then apply exploratory data analysis to the second subgroup to inform how we draw inferences about the first. Any outcome that shows a significant effect in both groups is defined as a “replicable” finding. (Bogomolov and Heller [2023]). In Section 5, we detail how we apply this non-random cross-screening method to detect replicable outcomes and test the global null hypothesis, i.e., identifying outcomes that are affected in at least one of the two subpopulations.
The remainder of the document is organized as follows. We introduce the BRFSS dataset and describe our data processing in Section 2. In Section 3, we report our candidate sets of combinations. In Section 4, we outline our stratification methodology. Finally, we conclude with a description of our data analysis procedure and explanation of several specifics of implementation in Section 5.
2. The Behavioral Risk Factor Surveillance System (BRFSS)
This study utilizes data from BRFSS, an ongoing health-related telephone survey system administered by the CDC. BRFSS collects comprehensive data on a wide range of health-related behaviors, chronic health conditions, healthcare access, and preventive service usage from the non-institutionalized adult population (18 years and older). The survey employs a dual-frame sampling methodology, using both landline and cellular telephone surveys to ensure representativeness of the data. The dataset collected in 2023 is the latest data available, has 433,323 observations from 48 states, the District of Columbia, Guam, Puerto Rico, and the US Virgin Islands, and includes topics such as health status, chronic conditions, health behaviors, healthcare access, and preventive services, as well as optional modules including social determinants of health, cancer screenings, and cognitive decline.
For the purpose of our study, we focus on the 10-state [^1] Delaware, Florida, Georgia, Missouri, Nevada, New Jersey, Oregon, Rhode Island, Tennessee, Virginia. sample that administered the CDC’s ACEs module in 2023. Following previous studies utilizing the BRFSS data (Bhan et al. [2014]; Gilbert et al. [2015]; Waehrer et al. [2020]), we collapse different questions into particular indicators. More details are provided in table 1. In total, we create 10 ACEs items from the BRFSS data (see table 2), which is consistent with the categorization in CDC-Kaiser study (Felitti et al. [1998]). It is worth noting that we first follow those criteria for data pre-processing, but these steps could be adjusted when we carry out exploratory data analysis in the data turnover..
Indicator Questions Collapsed Encoding Method
Any Sexual Abuse Number of times forced to touch, be touched, or have sex with anyone at least five years older or an adult Collapsed into a single indicator for any sexual abuse
Household Substance Use Alcohol and illegal drug use in the household Combined into a single indicator for household substance use
Physical Abuse Frequency of physical abuse Encoded as 1 if the answer is “once” or “more than once”
Verbal Abuse Frequency of verbal abuse Encoded as 1 if the answer is “once” or “more than once”
Parental Divorce Response to parental divorce question Affirmative response encoded as 1
Living with a Depressed or Mentally Ill Person Response to question about living with a mentally ill person Affirmative response encoded as 1
Living with an Incarcerated Person Response to question about living with an incarcerated person Affirmative response encoded as 1
Physical Violence Between Parents Frequency of physical violence between parents Affirmative response encoded as 1
Safety Did an adult make you feel safe and protected ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
Basic Needs Met Did an adult make sure basic needs were met ‘Never’ and ‘A little of the time’ encoded as 0, ‘Some of the time’, ‘Most of the time’, ‘All of the time’ encoded as 1
Table 1. Data Processing Steps and Collapsed Indicators
ACE Item Variable
Live With Anyone Depressed, Mentally Ill, Or Suicidal ACEDEPRS
Substance Abuse in the Home ACESUB
Live With Anyone Who Served Time in Prison or Jail ACEPRIS
Were Your Parents Divorced/Separated ACEDIVRC
How Often Did Your Parents Beat Each Other Up ACEPUNCH
How Often Did A Parent Physically Hurt You In Any Way ACEHURT1
How Often Did A Parent Swear At You ACESWEAR
Sexual Abuse ACESEX
Did An Adult Make You Feel Safe And Protected ACEADSAF
Did An Adult Make Sure Basic Needs Were Met ACEADNED
Table 2. ACE Items in BRFSS 2023
After processing the data, we restrict our sample to those who do not have any missing values in any ten ACEs variables. We also omit those individuals who do not have the appropriate measurement for depression. For any missing family background, we directly use the imputation value provided by the BRFSS 2023. In total, there are 49,811 remaining individuals, among whom 23,506 individuals are from blue states [^2] Delaware, New Jersey, Oregon, Rhode Island, Virginia (as established in the 2024 United States presidential election) and 26,305 individuals from red states [^3] Florida, Georgia, Missouri, Nevada, Tennessee. In the part of data with individuals from blue states, 8,076 individuals do not have any ACEs while 15,512 were exposed to at least one ACE. In the part of data with individuals from blue states, 9,163 individuals do not have any ACEs while 17,231 were exposed to at least one ACE.
3. Pre-specified Candidate Sets of ACEs
We collect ten variables of ACEs from the BRFSS 2023 dataset. To examine the diverse effects of combinations of ACEs on depression, we first construct candidate sets of ACEs that we are interested in. Since previous literature has identified four ACEs as a critical threshold, beyond which the effects appear to intensify significantly, and that individuals with at least four ACEs are at an increased risk for chronic disease, poor mental health, and negative social outcomes (Hughes et al. [2017]; Waehrer et al. [2020]), we consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs categories. We thus have 385 possible combinations of ACEs formed from ten candidate ACEs categories (see Appendix A).
4. Stratification
A previous study (Waehrer et al. [2020]) used BRFSS 2009 & 2012 to examine whether the relationship between ACEs and health outcomes, such as smoking, heavy drinking, and obesity, was similar across states, and controlled for baseline covariates including race/ethnicity, sex, and age. They also controlled for education, marital status and employment, which are post-treatment variables and likely to be affected by ACEs. Cox [1958] refers to these variables as “concomitant observations” and writes: “The essential point in our assumptions about these observations is that the value for any unit must be unaffected by the particular assignment of treatments to units actually used. In practice this means that either (a) the concomitant observations are taken before the assignment of treatments to units is made; or (b) the concomitant observations are made after the assignment of treatments, but before the effect of treatments has had time to develop…; or (c) we can assume from our knowledge of the nature of the concomitant observations concerned, that they are unaffected by the treatment.” If adjustments are made for variables that are affected by the treatment, then adjustments may remove part or all of the effect of the treatment, and hence estimators that adjust for a concomitant variable that has been affected by the treatment are generally biased (Rosenbaum [1984]). Therefore, we do not adjust for education, employment and marital status in our study.
Given the relatively small number of covariates, we seek to mitigate potential confounding bias by stratifying participants on their recorded baseline characteristics (race/ethnicity, age, sex) and then comparing those exposed to specific ACE combinations (the “treated” group) with those reporting no ACE exposure (the “control” group) within each stratum. Each hypothesis contrasts a particular ACE combination against having no ACEs. Individuals who have any additional ACEs beyond the combination of interest remain in the treated group. For example, if our focus is on the effect of substance abuse, participants exposed to both substance abuse and physical abuse are still classified under “substance abuse.” We conduct these stratifications separately for individuals residing in blue states and those in red states, resulting in $385\times 2$ distinct stratified analyses.
Specifically, we stratify participants by race/ethnicity, including White (non-Hispanic), Black (non-Hispanic), Hispanic, and Other (non-Hispanic); by age, with groups spanning 18-24, 25-34, 35-44, 45-54, 55-64, and 65+; and by sex, distinguishing between male and female participants. This results in a total of 48 distinct strata for both the blue and red subgroups. In the blue subgroup, the stratum with the lowest sample size contains 61 participants, with a median of 171. In the red subgroup, the stratum with the lowest sample size has 60 participants, with a median of 161. Detailed sample sizes for each stratum of both subgroups are provided in the appendix B.
5. Design of Data Analysis
Our inference will be based on stratified sets of participants in the BRFSS 2023, as described previously in Section 4. We aim to explore the effect of different combinations of ACEs in our aforementioned candidate sets on depression and also identify those with higher risk. We will deem those combinations of ACEs for which the effect of the treatment is found to be significantly increase the odds of depression in both subgroups - those residing in blue states and red states - as replicable findings. Also, we are interested in those combinations more than doubling the odds ratio of depression in only one subgroup.
5.1. Data Turnover Framework
We employ data turnover (Bekerman et al. in preparation, Bekerman et al. [2024a]) to analyze the data, ensuring control of the family-wise error rate (FWER; the chance of falsely rejecting any true null hypotheses at level $\alpha=0.05$) and facilitating flexible inference compared to automated cross-screening (Zhao et al. [2018]), by integrating both qualitative and quantitative insights from data exploration and formulating data-informed new hypotheses (e.g., Bekerman et al. [2024b]).
A brief description is as follows. We split our data into two subgroups: those individuals in our study residing in blue states and those residing in red states. The first subgroup, without data exploration, suggests a limited number of combinations of ACEs to be validated in the second subgroup based on some pre-determined selection procedure. Then, we do ad hoc data exploration on the second subgroup to choose how to conduct one-sided hypothesis tests on the first subgroup, possibly allowing for novel hypotheses to be identified and tested. In other words, both subgroups are used to screen and both subgroups are used to validate, but only one subgroup can generate novel hypotheses to be validated in the other subgroup while still maintaining the validity of our entire procedure. Because a limited number of combinations seek validation on the second subgroup, compared with directly controlling the FWER of 385 hypotheses on the whole data (eg. by using Bonferroni correction or Holm procedure), only a small fraction of the 385 potential hypotheses are considered in the validation sample.
5.2. Analysis Plan
We are interested in testing whether a combination of ACEs would increase the odds of depression compared with the reference population.
We plan to use the Cochran–Mantel–Haenszel test (CMH) in each subgroup to screen ACEs combinations. In the screening stage, we also need to define a cutoff of p-value (eg. 0.05). In a paper defining multimorbidity (Silber et al. [2018]), they chose form a 99.9% confidence interval in checking if the interval for the odds ratio was completely above 2. This is a stringent threshold to control for the number of hypotheses falsely passing the screening step. Yet, choosing this cutoff is rather heuristic given the sample sizes and number of hypotheses. A basic intuition would be: if the cutoff is too large, then almost all of the hypotheses would pass the screening stage, which would make this step meaningless; if the cutoff is too small, though only a few number of true nulls would falsely pass the screening stage, some true non-nulls would be screened out if the signals aren’t that strong.
Thus, following the structure of the BRFSS 2023, we set the number of participants in each subgroup at 1000, and set the number of hypotheses at 385, and try several cutoffs of p-value (0.001, 0.01,0.025, 0.05, 0.1), also varying the number of replicable hypotheses (1,50,100,150,200,250,300,350,385) and the signal-to-noise ratio (a sequence from 0.5 to 5, with a common difference 0.1). The result (see Figure 1) shows that whatever the number of replicable hypotheses and signal-to-noise ratio, the power using 0.001 as the cutoff is always the lowest. But there’s only trivial difference of power between choosing 0.01,0.025, 0.05 or 0.1, and none of them consistently perform the best under different values of the number of replicable hypotheses and the signal-to-noise ratio.
Figure 1. Simulation results of power with p-value cutoffs, number of replicable hypotheses and different signal-to-noise ratios.
Note: the $x$ axis represents the signal-to-noise ratio, the the $y$ axis represents the power. The facet titles in the gray bars are numbers of replicable hypotheses.
We intend to control the FWER at level $\alpha=0.05$. Therefore, we will apply Holm’s procedure (Holm [1979]) on each lists of p-values of selected hypotheses at level $\alpha/2=0.025$. Accordingly, we choose 0.025 as our cutoff in the screening stage. In other words, if the p-value of one hypothesis given by the CMH test is smaller than 0.025, then this hypothesis will pass the screening step, otherwise, this will be screened out. After selecting a list of hypotheses from subgroup 1 from the candidate sets collection, we will test them on subgroup 2. Specifically, for each selected hypothesis, we calculate the p-value using the data from subgroup 2, in the same way as we did in the screening stage on subgroup 1, and utilize Holm’s method at 0.025 to validate. Then, we will create a list of hypotheses from subgroup 2, including those selected from the candidate sets collection and generated by EDA, and validate on subgroup 1.
Moreover, we need to determine whether it is more beneficial to conduct EDA on the smaller or larger sample. Simulation results in Bekerman’s paper (in preparation) suggest that performing EDA on the smaller subgroup—in our case, participants living in blue states—is advantageous. We anticipate that incorporating data exploration within this subgroup will strengthen our study’s overall design by enabling data-informed decision-making and facilitating the emergence of new hypotheses. This flexibility may allow us to identify more complex hypotheses or yield unexpected benefits. For example, we may explore alternative ways of grouping ACE categories, refine our criteria for encoding ACE items, or investigate potential effect modification.
5.3. The high-risk group
After employing data turnover, we define the group of combinations for which the global null is rejected, that is, at least in one subgroup, as the high-risk group; in other words, the relevant combinations of ACEs will be identified as placing individuals at higher risk of depression. The group is formed as the union of validated combinations from subgroup 1 and validated combinations from subgroup 2, and simultaneously, we can tell what combinations are replicable and are only validated on one subgroup. If any pre-specified hypotheses passes four tests - passes the screening stage in the first subgroup and the validation in the second subgroup, and also passes the screening stage in the second subgroup and the validation in the first subgroup - they will be deemed as replicable. For new hypotheses formulated in the EDA process, if they are validated in the other subgroup, related combinations will be in the high-risk group.
We also provide a parsimonious list without duplicating any set whose components are already on the list as a smaller set. In other words, if a combination A in the collection is the subset of another combination B, we will remove B from the collection. For example, if both hypotheses regarding substance abuse and substance abuse + physical abuse are rejected, only substance abuse will be included in the parsimonious list. Previous literature has confirmed that ACEs can have a cumulative effect: each additional type of adversity elevates the probability of unfavorable outcomes (Felitti et al. [1998]; Chapman et al. [2004]). Therefore, given this monotonicity assumption, the validation of combination A implies that combination B increases the odds ratio of depression.
In terms of replicable findings, other than combinations simultaneously rejected in both subgroups, we also include the combination of non-duplicate components from two different combinations respectively rejected in each subgroup, if this combination has not been contained, based on the monotonicity assumption. For example, if substance abuse is rejected on subgroup 1, but substance abuse + physical abuse is not (which does not make sense in real-world scenario but may happen due to some statistical reasons), and only substance abuse + physical abuse is rejected on subgroup 2, then substance abuse + physical abuse will be deemed replicable.
The following tree diagrams (Figure 2, 3) makes things clear visually. We organize the ACE combinations in a tree structure, where each layer represents combinations of 1, 2, 3, or 4 ACE items, with each node corresponding to a unique combination. If a combination is rejected in the first subgroup, a yellow checkmark is placed at the top-left corner of the node, while a green checkmark appears at the bottom-right corner if the combination is rejected in the second subgroup. According to the monotonicity assumption, if a node is marked with either a yellow or green check, the entire subtree beneath that node inherits the same mark. Additionally, it’s important to note that identical combinations may appear in different subtrees within the same layer. If a combination node is marked in one subtree, all nodes with the same combination in other subtrees will be similarly marked. Ultimately, a combination node that is marked with checkmarks from both subgroups represents a replicable finding. In constructing the parsimonious list, we prioritize nodes from top to bottom, truncating any subtrees of replicable nodes as they are considered finalized, thus preventing redundant checks for combinations that have already been deemed replicable.
In Example 1 (Figure 2), only substance abuse is rejected in subgroup 1, and this node is marked with a yellow checkmark. In contrast, only the combination substance abuse + physical abuse is rejected in subgroup 2, with a green checkmark placed on this node. Moving to the second layer, all nodes in the subtree of substance abuse inherit the yellow checkmark. As a result, the combination substance abuse + physical abuse is marked with both checkmarks and is considered replicable. Simultaneously, the subtrees of substance abuse + physical abuse are truncated, meaning that any combinations containing substance abuse and physical abuse are no longer considered for further evaluation. In Example 2 (Figure 3), only sexual abuse is rejected in subgroup 1, while only physical abuse is rejected in subgroup 2. Following the rules of placing checkmarks, sexual abuse + verbal abuse has two checkmarks and is deemed replicable. Also, the subtrees of sexual abuse + verbal abuse are truncated.
Figure 2. Tree Diagram: Example 1
Figure 3. Tree Diagram: Example 2
Appendix A Candidate Sets of ACEs (variable names used)
We consider all quadruples of ACEs categories, all triples of ACEs categories, all pairs of ACEs categories, and all individual ACEs category and form this candidate sets of ACEs.
Table 3. Candidate sets of ACEs
Item.1 Item.2 Item.3 Item.4
1 ACESUB
2 ACEPRISN
3 ACEDIVRC
4 ACEPUNCH
5 ACEHURT1
6 ACESWEAR
7 ACESEX
8 ACEADSAF
9 ACEADNED
10 ACEDEPRS
11 ACESUB ACEPRISN
12 ACESUB ACEDIVRC
13 ACESUB ACEPUNCH
14 ACESUB ACEHURT1
15 ACESUB ACESWEAR
16 ACESUB ACESEX
17 ACESUB ACEADSAF
18 ACESUB ACEADNED
19 ACESUB ACEDEPRS
20 ACEPRISN ACEDIVRC
21 ACEPRISN ACEPUNCH
22 ACEPRISN ACEHURT1
23 ACEPRISN ACESWEAR
24 ACEPRISN ACESEX
25 ACEPRISN ACEADSAF
26 ACEPRISN ACEADNED
27 ACEPRISN ACEDEPRS
28 ACEDIVRC ACEPUNCH
29 ACEDIVRC ACEHURT1
30 ACEDIVRC ACESWEAR
31 ACEDIVRC ACESEX
32 ACEDIVRC ACEADSAF
33 ACEDIVRC ACEADNED
34 ACEDIVRC ACEDEPRS
35 ACEPUNCH ACEHURT1
36 ACEPUNCH ACESWEAR
37 ACEPUNCH ACESEX
38 ACEPUNCH ACEADSAF
39 ACEPUNCH ACEADNED
40 ACEPUNCH ACEDEPRS
41 ACEHURT1 ACESWEAR
42 ACEHURT1 ACESEX
43 ACEHURT1 ACEADSAF
44 ACEHURT1 ACEADNED
45 ACEHURT1 ACEDEPRS
46 ACESWEAR ACESEX
47 ACESWEAR ACEADSAF
48 ACESWEAR ACEADNED
49 ACESWEAR ACEDEPRS
50 ACESEX ACEADSAF
51 ACESEX ACEADNED
52 ACESEX ACEDEPRS
53 ACEADSAF ACEADNED
54 ACEADSAF ACEDEPRS
55 ACEADNED ACEDEPRS
56 ACESUB ACEPRISN ACEDIVRC
57 ACESUB ACEPRISN ACEPUNCH
58 ACESUB ACEPRISN ACEHURT1
59 ACESUB ACEPRISN ACESWEAR
60 ACESUB ACEPRISN ACESEX
61 ACESUB ACEPRISN ACEADSAF
62 ACESUB ACEPRISN ACEADNED
63 ACESUB ACEPRISN ACEDEPRS
64 ACESUB ACEDIVRC ACEPUNCH
65 ACESUB ACEDIVRC ACEHURT1
66 ACESUB ACEDIVRC ACESWEAR
67 ACESUB ACEDIVRC ACESEX
68 ACESUB ACEDIVRC ACEADSAF
69 ACESUB ACEDIVRC ACEADNED
70 ACESUB ACEDIVRC ACEDEPRS
71 ACESUB ACEPUNCH ACEHURT1
72 ACESUB ACEPUNCH ACESWEAR
73 ACESUB ACEPUNCH ACESEX
74 ACESUB ACEPUNCH ACEADSAF
75 ACESUB ACEPUNCH ACEADNED
76 ACESUB ACEPUNCH ACEDEPRS
77 ACESUB ACEHURT1 ACESWEAR
78 ACESUB ACEHURT1 ACESEX
79 ACESUB ACEHURT1 ACEADSAF
80 ACESUB ACEHURT1 ACEADNED
81 ACESUB ACEHURT1 ACEDEPRS
82 ACESUB ACESWEAR ACESEX
83 ACESUB ACESWEAR ACEADSAF
84 ACESUB ACESWEAR ACEADNED
85 ACESUB ACESWEAR ACEDEPRS
86 ACESUB ACESEX ACEADSAF
87 ACESUB ACESEX ACEADNED
88 ACESUB ACESEX ACEDEPRS
89 ACESUB ACEADSAF ACEADNED
90 ACESUB ACEADSAF ACEDEPRS
91 ACESUB ACEADNED ACEDEPRS
92 ACEPRISN ACEDIVRC ACEPUNCH
93 ACEPRISN ACEDIVRC ACEHURT1
94 ACEPRISN ACEDIVRC ACESWEAR
95 ACEPRISN ACEDIVRC ACESEX
96 ACEPRISN ACEDIVRC ACEADSAF
97 ACEPRISN ACEDIVRC ACEADNED
98 ACEPRISN ACEDIVRC ACEDEPRS
99 ACEPRISN ACEPUNCH ACEHURT1
100 ACEPRISN ACEPUNCH ACESWEAR
101 ACEPRISN ACEPUNCH ACESEX
102 ACEPRISN ACEPUNCH ACEADSAF
103 ACEPRISN ACEPUNCH ACEADNED
104 ACEPRISN ACEPUNCH ACEDEPRS
105 ACEPRISN ACEHURT1 ACESWEAR
106 ACEPRISN ACEHURT1 ACESEX
107 ACEPRISN ACEHURT1 ACEADSAF
108 ACEPRISN ACEHURT1 ACEADNED
109 ACEPRISN ACEHURT1 ACEDEPRS
110 ACEPRISN ACESWEAR ACESEX
111 ACEPRISN ACESWEAR ACEADSAF
112 ACEPRISN ACESWEAR ACEADNED
113 ACEPRISN ACESWEAR ACEDEPRS
114 ACEPRISN ACESEX ACEADSAF
115 ACEPRISN ACESEX ACEADNED
116 ACEPRISN ACESEX ACEDEPRS
117 ACEPRISN ACEADSAF ACEADNED
118 ACEPRISN ACEADSAF ACEDEPRS
119 ACEPRISN ACEADNED ACEDEPRS
120 ACEDIVRC ACEPUNCH ACEHURT1
121 ACEDIVRC ACEPUNCH ACESWEAR
122 ACEDIVRC ACEPUNCH ACESEX
123 ACEDIVRC ACEPUNCH ACEADSAF
124 ACEDIVRC ACEPUNCH ACEADNED
125 ACEDIVRC ACEPUNCH ACEDEPRS
126 ACEDIVRC ACEHURT1 ACESWEAR
127 ACEDIVRC ACEHURT1 ACESEX
128 ACEDIVRC ACEHURT1 ACEADSAF
129 ACEDIVRC ACEHURT1 ACEADNED
130 ACEDIVRC ACEHURT1 ACEDEPRS
131 ACEDIVRC ACESWEAR ACESEX
132 ACEDIVRC ACESWEAR ACEADSAF
133 ACEDIVRC ACESWEAR ACEADNED
134 ACEDIVRC ACESWEAR ACEDEPRS
135 ACEDIVRC ACESEX ACEADSAF
136 ACEDIVRC ACESEX ACEADNED
137 ACEDIVRC ACESEX ACEDEPRS
138 ACEDIVRC ACEADSAF ACEADNED
139 ACEDIVRC ACEADSAF ACEDEPRS
140 ACEDIVRC ACEADNED ACEDEPRS
141 ACEPUNCH ACEHURT1 ACESWEAR
142 ACEPUNCH ACEHURT1 ACESEX
143 ACEPUNCH ACEHURT1 ACEADSAF
144 ACEPUNCH ACEHURT1 ACEADNED
145 ACEPUNCH ACEHURT1 ACEDEPRS
146 ACEPUNCH ACESWEAR ACESEX
147 ACEPUNCH ACESWEAR ACEADSAF
148 ACEPUNCH ACESWEAR ACEADNED
149 ACEPUNCH ACESWEAR ACEDEPRS
150 ACEPUNCH ACESEX ACEADSAF
151 ACEPUNCH ACESEX ACEADNED
152 ACEPUNCH ACESEX ACEDEPRS
153 ACEPUNCH ACEADSAF ACEADNED
154 ACEPUNCH ACEADSAF ACEDEPRS
155 ACEPUNCH ACEADNED ACEDEPRS
156 ACEHURT1 ACESWEAR ACESEX
157 ACEHURT1 ACESWEAR ACEADSAF
158 ACEHURT1 ACESWEAR ACEADNED
159 ACEHURT1 ACESWEAR ACEDEPRS
160 ACEHURT1 ACESEX ACEADSAF
161 ACEHURT1 ACESEX ACEADNED
162 ACEHURT1 ACESEX ACEDEPRS
163 ACEHURT1 ACEADSAF ACEADNED
164 ACEHURT1 ACEADSAF ACEDEPRS
165 ACEHURT1 ACEADNED ACEDEPRS
166 ACESWEAR ACESEX ACEADSAF
167 ACESWEAR ACESEX ACEADNED
168 ACESWEAR ACESEX ACEDEPRS
169 ACESWEAR ACEADSAF ACEADNED
170 ACESWEAR ACEADSAF ACEDEPRS
171 ACESWEAR ACEADNED ACEDEPRS
172 ACESEX ACEADSAF ACEADNED
173 ACESEX ACEADSAF ACEDEPRS
174 ACESEX ACEADNED ACEDEPRS
175 ACEADSAF ACEADNED ACEDEPRS
176 ACESUB ACEPRISN ACEDIVRC ACEPUNCH
177 ACESUB ACEPRISN ACEDIVRC ACEHURT1
178 ACESUB ACEPRISN ACEDIVRC ACESWEAR
179 ACESUB ACEPRISN ACEDIVRC ACESEX
180 ACESUB ACEPRISN ACEDIVRC ACEADSAF
181 ACESUB ACEPRISN ACEDIVRC ACEADNED
182 ACESUB ACEPRISN ACEDIVRC ACEDEPRS
183 ACESUB ACEPRISN ACEPUNCH ACEHURT1
184 ACESUB ACEPRISN ACEPUNCH ACESWEAR
185 ACESUB ACEPRISN ACEPUNCH ACESEX
186 ACESUB ACEPRISN ACEPUNCH ACEADSAF
187 ACESUB ACEPRISN ACEPUNCH ACEADNED
188 ACESUB ACEPRISN ACEPUNCH ACEDEPRS
189 ACESUB ACEPRISN ACEHURT1 ACESWEAR
190 ACESUB ACEPRISN ACEHURT1 ACESEX
191 ACESUB ACEPRISN ACEHURT1 ACEADSAF
192 ACESUB ACEPRISN ACEHURT1 ACEADNED
193 ACESUB ACEPRISN ACEHURT1 ACEDEPRS
194 ACESUB ACEPRISN ACESWEAR ACESEX
195 ACESUB ACEPRISN ACESWEAR ACEADSAF
196 ACESUB ACEPRISN ACESWEAR ACEADNED
197 ACESUB ACEPRISN ACESWEAR ACEDEPRS
198 ACESUB ACEPRISN ACESEX ACEADSAF
199 ACESUB ACEPRISN ACESEX ACEADNED
200 ACESUB ACEPRISN ACESEX ACEDEPRS
201 ACESUB ACEPRISN ACEADSAF ACEADNED
202 ACESUB ACEPRISN ACEADSAF ACEDEPRS
203 ACESUB ACEPRISN ACEADNED ACEDEPRS
204 ACESUB ACEDIVRC ACEPUNCH ACEHURT1
205 ACESUB ACEDIVRC ACEPUNCH ACESWEAR
206 ACESUB ACEDIVRC ACEPUNCH ACESEX
207 ACESUB ACEDIVRC ACEPUNCH ACEADSAF
208 ACESUB ACEDIVRC ACEPUNCH ACEADNED
209 ACESUB ACEDIVRC ACEPUNCH ACEDEPRS
210 ACESUB ACEDIVRC ACEHURT1 ACESWEAR
211 ACESUB ACEDIVRC ACEHURT1 ACESEX
212 ACESUB ACEDIVRC ACEHURT1 ACEADSAF
213 ACESUB ACEDIVRC ACEHURT1 ACEADNED
214 ACESUB ACEDIVRC ACEHURT1 ACEDEPRS
215 ACESUB ACEDIVRC ACESWEAR ACESEX
216 ACESUB ACEDIVRC ACESWEAR ACEADSAF
217 ACESUB ACEDIVRC ACESWEAR ACEADNED
218 ACESUB ACEDIVRC ACESWEAR ACEDEPRS
219 ACESUB ACEDIVRC ACESEX ACEADSAF
220 ACESUB ACEDIVRC ACESEX ACEADNED
221 ACESUB ACEDIVRC ACESEX ACEDEPRS
222 ACESUB ACEDIVRC ACEADSAF ACEADNED
223 ACESUB ACEDIVRC ACEADSAF ACEDEPRS
224 ACESUB ACEDIVRC ACEADNED ACEDEPRS
225 ACESUB ACEPUNCH ACEHURT1 ACESWEAR
226 ACESUB ACEPUNCH ACEHURT1 ACESEX
227 ACESUB ACEPUNCH ACEHURT1 ACEADSAF
228 ACESUB ACEPUNCH ACEHURT1 ACEADNED
229 ACESUB ACEPUNCH ACEHURT1 ACEDEPRS
230 ACESUB ACEPUNCH ACESWEAR ACESEX
231 ACESUB ACEPUNCH ACESWEAR ACEADSAF
232 ACESUB ACEPUNCH ACESWEAR ACEADNED
233 ACESUB ACEPUNCH ACESWEAR ACEDEPRS
234 ACESUB ACEPUNCH ACESEX ACEADSAF
235 ACESUB ACEPUNCH ACESEX ACEADNED
236 ACESUB ACEPUNCH ACESEX ACEDEPRS
237 ACESUB ACEPUNCH ACEADSAF ACEADNED
238 ACESUB ACEPUNCH ACEADSAF ACEDEPRS
239 ACESUB ACEPUNCH ACEADNED ACEDEPRS
240 ACESUB ACEHURT1 ACESWEAR ACESEX
241 ACESUB ACEHURT1 ACESWEAR ACEADSAF
242 ACESUB ACEHURT1 ACESWEAR ACEADNED
243 ACESUB ACEHURT1 ACESWEAR ACEDEPRS
244 ACESUB ACEHURT1 ACESEX ACEADSAF
245 ACESUB ACEHURT1 ACESEX ACEADNED
246 ACESUB ACEHURT1 ACESEX ACEDEPRS
247 ACESUB ACEHURT1 ACEADSAF ACEADNED
248 ACESUB ACEHURT1 ACEADSAF ACEDEPRS
249 ACESUB ACEHURT1 ACEADNED ACEDEPRS
250 ACESUB ACESWEAR ACESEX ACEADSAF
251 ACESUB ACESWEAR ACESEX ACEADNED
252 ACESUB ACESWEAR ACESEX ACEDEPRS
253 ACESUB ACESWEAR ACEADSAF ACEADNED
254 ACESUB ACESWEAR ACEADSAF ACEDEPRS
255 ACESUB ACESWEAR ACEADNED ACEDEPRS
256 ACESUB ACESEX ACEADSAF ACEADNED
257 ACESUB ACESEX ACEADSAF ACEDEPRS
258 ACESUB ACESEX ACEADNED ACEDEPRS
259 ACESUB ACEADSAF ACEADNED ACEDEPRS
260 ACEPRISN ACEDIVRC ACEPUNCH ACEHURT1
261 ACEPRISN ACEDIVRC ACEPUNCH ACESWEAR
262 ACEPRISN ACEDIVRC ACEPUNCH ACESEX
263 ACEPRISN ACEDIVRC ACEPUNCH ACEADSAF
264 ACEPRISN ACEDIVRC ACEPUNCH ACEADNED
265 ACEPRISN ACEDIVRC ACEPUNCH ACEDEPRS
266 ACEPRISN ACEDIVRC ACEHURT1 ACESWEAR
267 ACEPRISN ACEDIVRC ACEHURT1 ACESEX
268 ACEPRISN ACEDIVRC ACEHURT1 ACEADSAF
269 ACEPRISN ACEDIVRC ACEHURT1 ACEADNED
270 ACEPRISN ACEDIVRC ACEHURT1 ACEDEPRS
271 ACEPRISN ACEDIVRC ACESWEAR ACESEX
272 ACEPRISN ACEDIVRC ACESWEAR ACEADSAF
273 ACEPRISN ACEDIVRC ACESWEAR ACEADNED
274 ACEPRISN ACEDIVRC ACESWEAR ACEDEPRS
275 ACEPRISN ACEDIVRC ACESEX ACEADSAF
276 ACEPRISN ACEDIVRC ACESEX ACEADNED
277 ACEPRISN ACEDIVRC ACESEX ACEDEPRS
278 ACEPRISN ACEDIVRC ACEADSAF ACEADNED
279 ACEPRISN ACEDIVRC ACEADSAF ACEDEPRS
280 ACEPRISN ACEDIVRC ACEADNED ACEDEPRS
281 ACEPRISN ACEPUNCH ACEHURT1 ACESWEAR
282 ACEPRISN ACEPUNCH ACEHURT1 ACESEX
283 ACEPRISN ACEPUNCH ACEHURT1 ACEADSAF
284 ACEPRISN ACEPUNCH ACEHURT1 ACEADNED
285 ACEPRISN ACEPUNCH ACEHURT1 ACEDEPRS
286 ACEPRISN ACEPUNCH ACESWEAR ACESEX
287 ACEPRISN ACEPUNCH ACESWEAR ACEADSAF
288 ACEPRISN ACEPUNCH ACESWEAR ACEADNED
289 ACEPRISN ACEPUNCH ACESWEAR ACEDEPRS
290 ACEPRISN ACEPUNCH ACESEX ACEADSAF
291 ACEPRISN ACEPUNCH ACESEX ACEADNED
292 ACEPRISN ACEPUNCH ACESEX ACEDEPRS
293 ACEPRISN ACEPUNCH ACEADSAF ACEADNED
294 ACEPRISN ACEPUNCH ACEADSAF ACEDEPRS
295 ACEPRISN ACEPUNCH ACEADNED ACEDEPRS
296 ACEPRISN ACEHURT1 ACESWEAR ACESEX
297 ACEPRISN ACEHURT1 ACESWEAR ACEADSAF
298 ACEPRISN ACEHURT1 ACESWEAR ACEADNED
299 ACEPRISN ACEHURT1 ACESWEAR ACEDEPRS
300 ACEPRISN ACEHURT1 ACESEX ACEADSAF
301 ACEPRISN ACEHURT1 ACESEX ACEADNED
302 ACEPRISN ACEHURT1 ACESEX ACEDEPRS
303 ACEPRISN ACEHURT1 ACEADSAF ACEADNED
304 ACEPRISN ACEHURT1 ACEADSAF ACEDEPRS
305 ACEPRISN ACEHURT1 ACEADNED ACEDEPRS
306 ACEPRISN ACESWEAR ACESEX ACEADSAF
307 ACEPRISN ACESWEAR ACESEX ACEADNED
308 ACEPRISN ACESWEAR ACESEX ACEDEPRS
309 ACEPRISN ACESWEAR ACEADSAF ACEADNED
310 ACEPRISN ACESWEAR ACEADSAF ACEDEPRS
311 ACEPRISN ACESWEAR ACEADNED ACEDEPRS
312 ACEPRISN ACESEX ACEADSAF ACEADNED
313 ACEPRISN ACESEX ACEADSAF ACEDEPRS
314 ACEPRISN ACESEX ACEADNED ACEDEPRS
315 ACEPRISN ACEADSAF ACEADNED ACEDEPRS
316 ACEDIVRC ACEPUNCH ACEHURT1 ACESWEAR
317 ACEDIVRC ACEPUNCH ACEHURT1 ACESEX
318 ACEDIVRC ACEPUNCH ACEHURT1 ACEADSAF
319 ACEDIVRC ACEPUNCH ACEHURT1 ACEADNED
320 ACEDIVRC ACEPUNCH ACEHURT1 ACEDEPRS
321 ACEDIVRC ACEPUNCH ACESWEAR ACESEX
322 ACEDIVRC ACEPUNCH ACESWEAR ACEADSAF
323 ACEDIVRC ACEPUNCH ACESWEAR ACEADNED
324 ACEDIVRC ACEPUNCH ACESWEAR ACEDEPRS
325 ACEDIVRC ACEPUNCH ACESEX ACEADSAF
326 ACEDIVRC ACEPUNCH ACESEX ACEADNED
327 ACEDIVRC ACEPUNCH ACESEX ACEDEPRS
328 ACEDIVRC ACEPUNCH ACEADSAF ACEADNED
329 ACEDIVRC ACEPUNCH ACEADSAF ACEDEPRS
330 ACEDIVRC ACEPUNCH ACEADNED ACEDEPRS
331 ACEDIVRC ACEHURT1 ACESWEAR ACESEX
332 ACEDIVRC ACEHURT1 ACESWEAR ACEADSAF
333 ACEDIVRC ACEHURT1 ACESWEAR ACEADNED
334 ACEDIVRC ACEHURT1 ACESWEAR ACEDEPRS
335 ACEDIVRC ACEHURT1 ACESEX ACEADSAF
336 ACEDIVRC ACEHURT1 ACESEX ACEADNED
337 ACEDIVRC ACEHURT1 ACESEX ACEDEPRS
338 ACEDIVRC ACEHURT1 ACEADSAF ACEADNED
339 ACEDIVRC ACEHURT1 ACEADSAF ACEDEPRS
340 ACEDIVRC ACEHURT1 ACEADNED ACEDEPRS
341 ACEDIVRC ACESWEAR ACESEX ACEADSAF
342 ACEDIVRC ACESWEAR ACESEX ACEADNED
343 ACEDIVRC ACESWEAR ACESEX ACEDEPRS
344 ACEDIVRC ACESWEAR ACEADSAF ACEADNED
345 ACEDIVRC ACESWEAR ACEADSAF ACEDEPRS
346 ACEDIVRC ACESWEAR ACEADNED ACEDEPRS
347 ACEDIVRC ACESEX ACEADSAF ACEADNED
348 ACEDIVRC ACESEX ACEADSAF ACEDEPRS
349 ACEDIVRC ACESEX ACEADNED ACEDEPRS
350 ACEDIVRC ACEADSAF ACEADNED ACEDEPRS
351 ACEPUNCH ACEHURT1 ACESWEAR ACESEX
352 ACEPUNCH ACEHURT1 ACESWEAR ACEADSAF
353 ACEPUNCH ACEHURT1 ACESWEAR ACEADNED
354 ACEPUNCH ACEHURT1 ACESWEAR ACEDEPRS
355 ACEPUNCH ACEHURT1 ACESEX ACEADSAF
356 ACEPUNCH ACEHURT1 ACESEX ACEADNED
357 ACEPUNCH ACEHURT1 ACESEX ACEDEPRS
358 ACEPUNCH ACEHURT1 ACEADSAF ACEADNED
359 ACEPUNCH ACEHURT1 ACEADSAF ACEDEPRS
360 ACEPUNCH ACEHURT1 ACEADNED ACEDEPRS
361 ACEPUNCH ACESWEAR ACESEX ACEADSAF
362 ACEPUNCH ACESWEAR ACESEX ACEADNED
363 ACEPUNCH ACESWEAR ACESEX ACEDEPRS
364 ACEPUNCH ACESWEAR ACEADSAF ACEADNED
365 ACEPUNCH ACESWEAR ACEADSAF ACEDEPRS
366 ACEPUNCH ACESWEAR ACEADNED ACEDEPRS
367 ACEPUNCH ACESEX ACEADSAF ACEADNED
368 ACEPUNCH ACESEX ACEADSAF ACEDEPRS
369 ACEPUNCH ACESEX ACEADNED ACEDEPRS
370 ACEPUNCH ACEADSAF ACEADNED ACEDEPRS
371 ACEHURT1 ACESWEAR ACESEX ACEADSAF
372 ACEHURT1 ACESWEAR ACESEX ACEADNED
373 ACEHURT1 ACESWEAR ACESEX ACEDEPRS
374 ACEHURT1 ACESWEAR ACEADSAF ACEADNED
375 ACEHURT1 ACESWEAR ACEADSAF ACEDEPRS
376 ACEHURT1 ACESWEAR ACEADNED ACEDEPRS
377 ACEHURT1 ACESEX ACEADSAF ACEADNED
378 ACEHURT1 ACESEX ACEADSAF ACEDEPRS
379 ACEHURT1 ACESEX ACEADNED ACEDEPRS
380 ACEHURT1 ACEADSAF ACEADNED ACEDEPRS
381 ACESWEAR ACESEX ACEADSAF ACEADNED
382 ACESWEAR ACESEX ACEADSAF ACEDEPRS
383 ACESWEAR ACESEX ACEADNED ACEDEPRS
384 ACESWEAR ACEADSAF ACEADNED ACEDEPRS
385 ACESEX ACEADSAF ACEADNED ACEDEPRS
Appendix B Sample Sizes by Stratum in Each Subgroup
Table 4. Sample Sizes by Stratum in the Blue Subgroup
Race/Ethnicity Sex Age Sample Size
1 White, Non-Hispanic Male 18-24 362
2 Black, Non-Hispanic Male 18-24 61
3 Hispanic Male 18-24 182
4 Other race, Non-Hispanic Male 18-24 91
5 White, Non-Hispanic Female 18-24 232
6 Black, Non-Hispanic Female 18-24 63
7 Hispanic Female 18-24 135
8 Other race, Non-Hispanic Female 18-24 72
9 White, Non-Hispanic Male 25-34 695
10 Black, Non-Hispanic Male 25-34 82
11 Hispanic Male 25-34 234
12 Other race, Non-Hispanic Male 25-34 133
13 White, Non-Hispanic Female 25-34 621
14 Black, Non-Hispanic Female 25-34 99
15 Hispanic Female 25-34 274
16 Other race, Non-Hispanic Female 25-34 125
17 White, Non-Hispanic Male 35-44 820
18 Black, Non-Hispanic Male 35-44 103
19 Hispanic Male 35-44 222
20 Other race, Non-Hispanic Male 35-44 194
21 White, Non-Hispanic Female 35-44 908
22 Black, Non-Hispanic Female 35-44 157
23 Hispanic Female 35-44 247
24 Other race, Non-Hispanic Female 35-44 129
25 White, Non-Hispanic Male 45-54 955
26 Black, Non-Hispanic Male 45-54 135
27 Hispanic Male 45-54 171
28 Other race, Non-Hispanic Male 45-54 131
29 White, Non-Hispanic Female 45-54 1212
30 Black, Non-Hispanic Female 45-54 168
31 Hispanic Female 45-54 206
32 Other race, Non-Hispanic Female 45-54 151
33 White, Non-Hispanic Male 55-64 1513
34 Black, Non-Hispanic Male 55-64 141
35 Hispanic Male 55-64 160
36 Other race, Non-Hispanic Male 55-64 132
37 White, Non-Hispanic Female 55-64 2054
38 Black, Non-Hispanic Female 55-64 267
39 Hispanic Female 55-64 171
40 Other race, Non-Hispanic Female 55-64 110
41 White, Non-Hispanic Male 65+ 3688
42 Black, Non-Hispanic Male 65+ 260
43 Hispanic Male 65+ 118
44 Other race, Non-Hispanic Male 65+ 180
45 White, Non-Hispanic Female 65+ 4561
46 Black, Non-Hispanic Female 65+ 473
47 Hispanic Female 65+ 139
48 Other race, Non-Hispanic Female 65+ 169
Table 5. Sample Sizes by Stratum in the Red Subgroup
Race/Ethnicity Sex Age Sample Size
1 White, Non-Hispanic Male 18-24 362
2 Black, Non-Hispanic Male 18-24 61
3 Hispanic Male 18-24 182
4 Other race, Non-Hispanic Male 18-24 91
5 White, Non-Hispanic Female 18-24 232
6 Black, Non-Hispanic Female 18-24 63
7 Hispanic Female 18-24 135
8 Other race, Non-Hispanic Female 18-24 72
9 White, Non-Hispanic Male 25-34 695
10 Black, Non-Hispanic Male 25-34 82
11 Hispanic Male 25-34 234
12 Other race, Non-Hispanic Male 25-34 133
13 White, Non-Hispanic Female 25-34 621
14 Black, Non-Hispanic Female 25-34 99
15 Hispanic Female 25-34 274
16 Other race, Non-Hispanic Female 25-34 125
17 White, Non-Hispanic Male 35-44 820
18 Black, Non-Hispanic Male 35-44 103
19 Hispanic Male 35-44 222
20 Other race, Non-Hispanic Male 35-44 194
21 White, Non-Hispanic Female 35-44 908
22 Black, Non-Hispanic Female 35-44 157
23 Hispanic Female 35-44 247
24 Other race, Non-Hispanic Female 35-44 129
25 White, Non-Hispanic Male 45-54 955
26 Black, Non-Hispanic Male 45-54 135
27 Hispanic Male 45-54 171
28 Other race, Non-Hispanic Male 45-54 131
29 White, Non-Hispanic Female 45-54 1212
30 Black, Non-Hispanic Female 45-54 168
31 Hispanic Female 45-54 206
32 Other race, Non-Hispanic Female 45-54 151
33 White, Non-Hispanic Male 55-64 1513
34 Black, Non-Hispanic Male 55-64 141
35 Hispanic Male 55-64 160
36 Other race, Non-Hispanic Male 55-64 132
37 White, Non-Hispanic Female 55-64 2054
38 Black, Non-Hispanic Female 55-64 267
39 Hispanic Female 55-64 171
40 Other race, Non-Hispanic Female 55-64 110
41 White, Non-Hispanic Male 65+ 3688
42 Black, Non-Hispanic Male 65+ 260
43 Hispanic Male 65+ 118
44 Other race, Non-Hispanic Male 65+ 180
45 White, Non-Hispanic Female 65+ 4561
46 Black, Non-Hispanic Female 65+ 473
47 Hispanic Female 65+ 139
48 Other race, Non-Hispanic Female 65+ 169